# Evan Munro    Xu Kuang    Stefan Wager

### v3

@@ B.6 Proof of Theorem 8
$\hat{\sigma}^{*2}_{D}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}\left(Y_{i}-y(1,\,P_{n})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}-y(0,\,P_{n})\right)}{1-\pi}-\Delta_{i}(W_{i},\,P_{n})\right)^{2}.$ (57)
@@
$\begin{split}&\hat{\sigma}^{*2}_{D}=n\operatorname{Var}_{T}\left[\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\left(Y_{i}-y(1,\,P_{n})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}-y(0,\,P_{n})\right)}{1-\pi}-\Delta_{i}(W_{i},\,P_{n})\right)\right],\end{split}$ (58)
where $\operatorname{Var}_{T}\left[.\right]$ denotes variance induced by only the random weights $T_{i}$. Now, by asymptotic equicontinuity as shown in Lemma 20,
$\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\left(Y_{i}-y(1,\,P_{n})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}-y(0,\,P_{n})\right)}{1-\pi}-\Delta_{i}(W_{i},\,P_{n})\right)$
$\quad\quad=\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\left(Y_{i}(1,\,p^{*}_{\pi})-y(1,\,p^{*}_{\pi})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}(0,\,p^{*}_{\pi})-y(0,\,p^{*}_{\pi})\right)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi})\right)$
$\quad\quad\quad\quad\quad\quad+\nabla_{p}^{\top}[y(1,\,p^{*}_{\pi})-y(0,\,p^{*}_{\pi})]\xi_{z}^{-1}\left(z\left(P_{n}\right)-z\left(p^{*}_{\pi}\right)\right)+B_{n}(\mathbf{T}),$
with $\lim_{n\rightarrow\infty}n\mathbb{E}\left[B_{n}(\mathbf{T})^{2}\right]=0$. Thus, from (58), we immediately see that
$\begin{split}\hat{\sigma}^{*2}_{D}&=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi})\right)^{2}+n\operatorname{Var}_{T}\left[B_{n}(\mathbf{T})\right]\\ &\quad\quad\quad\quad+2\operatorname{Cov}_{T}\left[\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi})\right),\,B_{n}(\mathbf{T})\right],\end{split}$ (59)
with notation $\varepsilon_{i}(w)$ as defined in the statement of Theorem 5. Now,
@@
and so the terms depending on $B_{n}(\mathbf{T})$ are asymptotically negligible; the covariance term needs to be bounded by Cauchy-Schwarz before applying (60). Finally, the dominant term in (59) is an IID average whose mean is $\sigma^{2}_{D}$, and thus our result follows by standard concentration arguments.
The consistency of $\hat{\sigma}_{I}^{2}$ can be established via an analogous argument, which is omitted here for conciseness.
@@ B.7 Proof of Proposition 9
$=\mathbb{E}\left[Y_{i}(w_{i}=1;W_{-k})-Y_{i}(w_{i}=0;W_{-i})|X_{i}=x\right]$
@@ B.8 Proof of Theorem 10
$\lim_{n\rightarrow\infty}\tau_{\text{CADE}}(x)-\mathbb{E}\left[Y_{i}\left(1,\,p^{*}_{\pi}\right)-Y_{i}\left(0,\,p^{*}_{\pi}\right)\,\big{|}\,X_{i}=x\right]=0,$
@@
For the second claim, we follow the structure from the proof of Theorem 2 and note that, by Proposition 9, it suffices to show that
$\mathbb{E}\left[\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\,\big{|}\,X_{k}\in S\right]=\mathbb{E}\left[\tau^{*}_{\text{CADE}}(x)+\tau^{*}_{\text{CAIE}}(x)\,\big{|}\,X_{k}\in S\right].$ (61)
We have chosen the set $S$ such that $\mathbb{P}\left[X_{k}\in S\right]>0$. Thus, there is a $c>0$ such that with probability $1-e^{cn}$ there is at least one observation $X_{k}\in S$, and so, using a convention that $/0=0$ 0,
$\mathbb{E}\left[\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\,\big{|}\,X_{k}\in S\right]=\mathbb{E}\left[\frac{1}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\right]+\mathcal{O}\left(e^{-cn}\right).$
@@
$\begin{split}&\mathbb{E}\left[\frac{1}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\right]\\ &\quad\quad\quad=\mathbb{E}\left[\tau^{*}_{\text{CADE}}(x)+\tau^{*}_{\text{CAIE}}(x)\,\big{|}\,X_{k}\in S\right]\mathbb{P}\left[\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert>0\right]\\ &\quad\quad\quad\quad\quad\quad+\mathbb{E}\left[\frac{n}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}B_{n}\right],\end{split}$ (62)
with $\lim_{n\rightarrow\infty}n\mathbb{E}\left[B_{n}^{2}\right]=0$. We have that
$\mathbb{E}\left[\left\lvert\sum_{X_{k}\in S}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}B_{n}\right\rvert\right]=0$
by Cauchy-Schwarz as in the proof of Lemma 15, and that the term inside the expectation admits an almost sure bound that scales polynomially in $n$. Meanwhile, $n\,/\,\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert\leq 2/\mathbb{P}\left[X_{k}\in S\right]$ with probability that decays exponentially in $n$. Thus, the whole expectation in (62) goes to 0 with $n$, as desired.
@@ B.9 Proof of Theorem 11
By treatment randomization |$\left\lvert\left\{i\in N_{k}(x):W_{i}=1\right\}\right\rvert/k\rightarrow_{p}\pi$, and so
$\hat{\tau}_{\text{CADE}}(x)=\frac{1}{k}\sum_{\left\{i\in N_{k}(x)\right\}}\left(\frac{W_{i}}{\pi}\,Y_{i}-\frac{1-W_{i}}{1-\pi}\,Y_{i}\right).$ (63)
Now, let $B_{\varepsilon}(x)=\left\{x^{\prime}:\left\lVert x-x^{\prime}\right\rVert_{2}\leq\varepsilon\right\}$. For any $\varepsilon>0$, we have $\mathbb{P}\left[X\in B_{\varepsilon}(x)\right]>0$ and so, because $k/n\rightarrow 0$, the $k$-nearest neighbor set $N_{k}(x)$ is asymptotically concentrated near $x$, i.e., $\lim_{n\rightarrow\infty}\mathbb{P}\left[\left\{X_{i}:i\in N_{k}(x)\right\}\subset B_{\varepsilon}(x)\right]=1$ for any $\varepsilon>0$. Thus, because the conditional distribution of $Y_{i}(w,\,p)$ given $X_{i}=x$ varies continuously in $x$ and $Y_{i}(w,\,p)$ is bounded, conditional moments of $Y_{i}(w,\,p)$ also vary continuously in $x$ and so we can follow the proof of Lemma 20 to verify that, for any sequence $p_{n}\rightarrow_{p}p$,
$\begin{split}\lim_{n\rightarrow\infty}\left\lvert\frac{1}{k}\sum_{N_{k}(x)}\frac{W_{i}}{\pi}\,Y_{i}(1,p_{n})-\frac{1}{k}\sum_{N_{k}(x)}\left(\frac{W_{i}}{\pi}\,Y_{i}(1,p)+y(1,p_{n};\,X_{i})-y(1,p;\,X_{i})\right)\right\rvert=o_{p}\left(\frac{1}{\sqrt{k}}\right),\end{split}$ (64)
and a similar expansion holds for $(1-W_{i})Y_{i}(0,\,p_{n})/(1-\pi)$. Because $P_{n}\rightarrow p^{*}_{\pi}$ by Theorem 1 and $y(w,\,p;\,x)$ is continuous in $p$ and $x$ (and thus uniformly continuous in $p$ in a neighborhood of $x$), we have that
$\frac{1}{k}\sum_{\left\{i\in N_{k}(x)\right\}}\left(\frac{W_{i}}{\pi}\,Y_{i}-\frac{1-W_{i}}{1-\pi}\,Y_{i}\right)=\frac{1}{k}\sum_{\left\{i\in N_{k}(x)\right\}}\left(\frac{W_{i}}{\pi}\,Y_{i}(1,\,p^{*}_{\pi})-\frac{1-W_{i}}{1-\pi}\,Y_{i}(0,\,p^{*}_{\pi})\right)+o_{p}(1).$ (65)
The result then follows by concentration because $k\rightarrow\infty$. We note that this argument can be extended to obtain rates of convergence under smoothness assumptions following Stone (1980).

### v4

@@ B.6 Proof of Theorem 8
$\hat{\sigma}^{*2}_{D}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}\left(Y_{i}-y_{n}(1,\,P_{n})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}-y_{n}(0,\,P_{n})\right)}{1-\pi}-\Delta_{i}(W_{i},\,P_{n}+U_{i})\right)^{2}.$ (62)
@@
$\begin{split}&\hat{\sigma}^{*2}_{D}=n\operatorname{Var}_{T}\left[\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\left(Y_{i}-y_{n}(1,\,P_{n})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}-y_{n}(0,\,P_{n})\right)}{1-\pi}-\Delta_{i}(W_{i},\,P_{n}+U_{i})\right)\right],\end{split}$ (63)
where $\operatorname{Var}_{T}\left[.\right]$ denotes variance induced by only the random weights $T_{i}$. Let
$F_{in}(p)=T_{i}\left(\frac{W_{i}\left(Y_{i}-y_{n}(1,\,P_{n})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}-y_{n}(0,\,P_{n})\right)}{1-\pi}-\Delta_{i}(W_{i},\,P_{n}+U_{i})\right).$
This is the sum of bounded functions that meet the conditions of Lemma 22, multiplied by a bounded and independent random variable, so the conclusions of Lemma 22 apply also to sample averages of $F_{in}(p)$. So, we can write
$\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\left(Y_{i}-y_{n}(1,\,P_{n})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}-y_{n}(0,\,P_{n})\right)}{1-\pi}-\Delta_{i}(W_{i},\,P_{n})\right)$
$\quad\quad=\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\left(Y_{i}(1,\,p^{*}_{\pi}+U_{i})-y_{n}(1,\,p^{*}_{\pi})\right)}{\pi}-\frac{(1-W_{i})\left(Y_{i}(0,\,p^{*}_{\pi}+U_{i})-y_{n}(0,\,p^{*}_{\pi})\right)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi}+U_{i})\right)$
$\quad\quad\quad\quad\quad\quad+\nabla_{p}^{\top}[y(1,\,p^{*}_{\pi})-y(0,\,p^{*}_{\pi})]\xi_{z}^{-1}\left(z_{n}\left(P_{n}\right)-z_{n}\left(p^{*}_{\pi}\right)\right)+B_{n}(\mathbf{T}),$
where $B_{n}(\mathbf{T})$ is the residual term $R_{n}$ from Lemma 22, making clear the dependence on $T_{i}$, so $\lim_{n\rightarrow\infty}n\mathbb{E}\left[B_{n}(\mathbf{T})^{2}\right]=0$. Thus, from (63), we immediately see that
$\begin{split}\hat{\sigma}^{*2}_{D}&=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}\varepsilon_{in}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{in}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi}+U_{i})\right)^{2}+n\operatorname{Var}_{T}\left[B_{n}(\mathbf{T})\right]\\ &\quad\quad\quad\quad+2\operatorname{Cov}_{T}\left[\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\varepsilon_{in}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{in}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi}+U_{i})\right),\,B_{n}(\mathbf{T})\right],\end{split}$ (64)
with notation $\varepsilon_{in}(w)=Y_{i}(w,p^{*}_{\pi}+U_{i})-y_{n}(w,p^{*}_{\pi})$, augmented from the definition of $\varepsilon_{i}(w)$ in the statement of Theorem 5. Now,
@@
and so the terms depending on $B_{n}(\mathbf{T})$ are asymptotically negligible; the covariance term needs to be bounded by Cauchy-Schwarz before applying (65), as below:
$n\operatorname{Cov}_{T}\left[\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi})\right),\,B_{n}(\mathbf{T})\right]$
$\quad\quad\leq\sqrt{n\operatorname{Var}_{T}\left[\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi})\right)\right]}\sqrt{n\mathbb{E}\left[\operatorname{Var}_{T}\left[B_{n}(\mathbf{T})\right]\right]}$
$\quad\quad=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}\varepsilon_{i}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{i}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi})\right)^{2}}\sqrt{o_{p}(1)}$
$\quad\quad=o_{p}(1).$
Finally, the dominant term in (64) is an IID average whose mean is
$\mathbb{E}\left[\left(\frac{W_{i}\varepsilon_{in}(1)}{\pi}-\frac{(1-W_{i})\varepsilon_{in}(0)}{1-\pi}-\Delta_{i}(W_{i},\,p^{*}_{\pi}+U_{i})\right)^{2}\right]=\sigma^{2}_{D}+o(1).$
The consistency of $\hat{\sigma}_{I}^{2}$ can be established via an analogous argument, which we omit for brevity.
@@ B.7 Proof of Proposition 9
$=\mathbb{E}\left[Y_{i}(w_{i}=1;W_{-i})-Y_{i}(w_{i}=0;W_{-i})|X_{i}=x\right]$
@@ B.8 Proof of Theorem 10
$\lim_{n\rightarrow\infty}\bar{\tau}_{\text{CADE}}(x)-\mathbb{E}\left[Y_{i}\left(1,\,p^{*}_{\pi}\right)-Y_{i}\left(0,\,p^{*}_{\pi}\right)\,\big{|}\,X_{i}=x\right]=0,$
@@
For the second claim, we follow the structure from the proof of Theorem 2 and note that, by Proposition 9, it suffices to show that, for any set $S$ such that $\mathbb{P}\left[X_{k}\in S\right]>0$,
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\,\big{|}\,X_{k}\in S\right]=\mathbb{E}\left[\tau^{*}_{\text{CADE}}(X_{k})+\tau^{*}_{\text{CAIE}}(X_{k})\,\big{|}\,X_{k}\in S\right].$ (66)
Now, fix such an $S$. Given that $\mathbb{P}\left[X_{k}\in S\right]>0$, there is a $c>0$ such that with probability $1-e^{cn}$ there is at least one observation $X_{k}\in S$, and so, using a convention that $/0=0$ 0,
$\mathbb{E}\left[\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\,\big{|}\,X_{k}\in S\right]=\mathbb{E}\left[\frac{1}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\right]+\mathcal{O}\left(e^{-cn}\right).$ (67)
@@
$\begin{split}&\mathbb{E}\left[\frac{1}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{\partial}{\partial\pi_{k}}\mathbb{E}_{\pi}\left[\sum_{i=1}^{n}Y_{i}\right]\right]\\ &\quad\quad\quad=\mathbb{E}\left[\tau^{*}_{\text{CADE}}(X_{k})+\tau^{*}_{\text{CAIE}}(X_{k})\,\big{|}\,X_{k}\in S\right]\mathbb{P}\left[\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert>0\right]\\ &\quad\quad\quad\quad\quad\quad+\mathbb{E}\left[\frac{n}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}B_{n}\right],\end{split}$ (68)
where the residual term $B_{n}$ is as defined in (57). Now, we already know from the proof of Lemma 15 that $\lim_{n\rightarrow\infty}n\mathbb{E}\left[B_{n}^{2}\right]=0$. By Cauchy-Schwarz,
$\mathbb{E}\left[\left\lvert\frac{n}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}B_{n}\right\rvert\right]\leq\sqrt{n\mathbb{E}\left[B_{n}^{2}\right]}\sqrt{n\mathbb{E}\left[\left(\frac{1}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}\right)^{2}\right]}.$
Because $0<\eta\leq\pi_{k}\leq 1-\eta$ in Design 1, we have
$n\mathbb{E}\left[\left(\frac{1}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}\right)^{2}\right]\leq\frac{n}{\eta(1-\eta)}\,\mathbb{E}\left[\frac{1\left(\left\{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert>0\right\}\right)}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\right],$
and the right-hand side term above converges to $1/\left(\eta(1-\eta)\mathbb{P}\left[X_{k}\in S\right]\right)$ as $n$ gets large. Thus,
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\frac{n}{\left\lvert\left\{k:X_{k}\in S\right\}\right\rvert}\sum_{X_{k}\in S}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}B_{n}\right]=0,$
and so (67) and (68) together imply (66).
@@ B.9 Proof of Theorem 11
For simplicity, we will assume throughout this proof that $X_{i}$ has a continuous distribution, and so the set of $k$-nearest neighbors is almost surely unique. [^12] If $X_{i}$ has a discrete distribution, we also need to address random tie-breaking; this introduces some extra notational overhead, but no conceptual difficulties. Now, by treatment randomization |$\left\lvert\left\{i\in N_{k}(x):W_{i}=1\right\}\right\rvert/k\rightarrow_{p}\pi$, and so because the outcomes are bounded
$\hat{\tau}_{\text{CADE}}(x)=\frac{1}{k}\sum_{\left\{i\in N_{k}(x)\right\}}\left(\frac{W_{i}}{\pi}\,Y_{i}-\frac{1-W_{i}}{1-\pi}\,Y_{i}\right)+o_{p}(1).$ (69)
Next, let $B_{\varepsilon}(x)=\left\{x^{\prime}:\left\lVert x-x^{\prime}\right\rVert_{2}\leq\varepsilon\right\}$, and define $\varepsilon_{n}(x)=\inf\left\{\varepsilon:\mathbb{P}\left[X_{i}\in B_{\varepsilon}(x)\right]\geq k/n\right\}$. When $X_{i}$ has a continuous distribution, we have $\mathbb{P}\left[\left\lVert X_{i}-x\right\rVert_{2}=\varepsilon_{n}(x)\right]=0$, and so
|$\left\lvert\left\{i:X_{i}\in B_{\varepsilon_{n}}(x)\right\}\right\rvert\sim\text{Binomial}(k/n,\,n).$
It immediately follows that the sets $\left\{i:X_{i}\in B_{\varepsilon_{n}}(x)\right\}$ and $N_{k}(x)$ disagree on only $\mathcal{O}_{p}(\sqrt{k})$ elements, and so (69) implies that (again because the $Y_{i}$ are bounded)
$\hat{\tau}_{\text{CADE}}(x)=\frac{1}{\left\lvert\left\{i:X_{i}\in B_{\varepsilon_{n}}(x)\right\}\right\rvert}\sum_{\left\{i:X_{i}\in B_{\varepsilon_{n}}(x)\right\}}\left(\frac{W_{i}}{\pi}\,Y_{i}-\frac{1-W_{i}}{1-\pi}\,Y_{i}\right)+o_{p}(1).$ (70)
We note that the leading term in the above expression is an IID average for any given $n$. Furthermore, because $\mathbb{P}\left[X\in B_{\varepsilon}(x)\right]>0$ for any $\varepsilon>0$ by assumption, we have $\varepsilon_{n}\rightarrow 0$; and, because the conditional distribution of $Y_{i}(w,\,p)$ given $X_{i}=x$ varies continuously in $x$ and $Y_{i}(w,\,p)$ is bounded, conditional moments of $Y_{i}(w,\,p)$ also vary continuously in $x$. Thus, the condition (89) from Lemma 22 holds for the potential outcome functions used to construct (70), and so for any sequence $p_{n}\rightarrow_{p}p$,
$\begin{split}&\lim_{n\rightarrow\infty}\Bigg{|}\frac{1}{\left\lvert\left\{i:X_{i}\in B_{\varepsilon_{n}}(x)\right\}\right\rvert}\sum_{\left\{i:X_{i}\in B_{\varepsilon_{n}}(x)\right\}}\bigg{(}\frac{W_{i}}{\pi}\,Y_{i}(1,p_{n})\\ &\quad\quad\quad\quad-\left(\frac{W_{i}}{\pi}\,Y_{i}(1,p)+y(1,p_{n};\,X_{i})-y(1,p;\,X_{i})\right)\bigg{)}\Bigg{|}=o_{p}\left(\frac{1}{\sqrt{k}}\right),\end{split}$ (71)
and a similar expansion holds for $(1-W_{i})Y_{i}(0,\,p_{n})/(1-\pi)$. Finally, because $P_{n}\rightarrow_{p}p^{*}_{\pi}$ by Theorem 1, $k\to\infty$, and $y(w,\,p;\,x)$ is continuous in $p$ and $x$ (and thus uniformly continuous in $p$ in a neighborhood of $x$), we can combine all above results to get
$\hat{\tau}_{\text{CADE}}(x)=\frac{1}{k}\sum_{\left\{i\in N_{k}(x)\right\}}\left(\frac{W_{i}}{\pi}\,Y_{i}(1,\,p^{*}_{\pi})-\frac{1-W_{i}}{1-\pi}\,Y_{i}(0,\,p^{*}_{\pi})\right)+o_{p}(1).$ (72)
We also note that this argument can be extended to obtain rates of convergence under smoothness assumptions following Stone (1980).